Can you please help with this word problem? Thanks

The period T (time in seconds for one complete
cycle) of a simple pendulum is related to the length L
(in feet) of the pendulum by the formula 8T^2= pi^2L. If a
child is on a swing with a 10-foot chain, then how long
does it take to complete one cycle of the swing?

Answer by Alan3354(21605)

The period T (time in seconds for one complete
cycle) of a simple pendulum is related to the length L
(in feet) of the pendulum by the formula 8T^2= pi^2L. If a
child is on a swing with a 10-foot chain, then how long
does it take to complete one cycle of the swing?
-----------------------
8T^2= pi^2L
L = 10
8T^2= 10*pi^2
T^2 = 5pi^2/4 seconds
T^2 = 1.25PI^2
T^2 = 12.337
T = 3.512 seconds